OpenStudy (onepieceftw):

Simplify: (sin Θ − cos Θ)^2 + (sin Θ + cos Θ)^2 I'm pretty sure it's two?

OpenStudy (xapproachesinfinity):

yes it is two can you show how though?

OpenStudy (triciaal):

(a-b)^2 + (a + b)^2 (a-b+ a + b)^2 - 2 (a-b)(a + b) (2a)^2 - 2 ( a^2 - b^2) 2a^2 + 2b^2 2(a^2 + b^2) a^2 + b^2 = 1 so 2(a^2 + b^2) = 2

OpenStudy (onepieceftw):

I squared the whole thing, then subtracted cos from cos and added sin to sin

OpenStudy (triciaal):

a = sin theta and b = cos theta

OpenStudy (onepieceftw):

pardon me, I took the square root of the whole thing.

OpenStudy (xapproachesinfinity):

your wording is vague, but you got the right answer the correct wording is: expand the squares, collect the terms and add up sin^2+con^2=1 -2cossin cancels with 2cossin

OpenStudy (xapproachesinfinity):

i don't see anywhere to do square root? that would be wrong

OpenStudy (onepieceftw):

Since both expression in the parenthesis are squared, it can be simplified by taking the square root

OpenStudy (xapproachesinfinity):

noo wrong! you can't use square root just because they squared

OpenStudy (xapproachesinfinity):

\[(\sin \theta -\cos \theta)^2+(\sin \theta +\cos \theta)^2 \] \[\sin^2\theta -2\sin \theta \cos \theta +\cos^2 \theta +\sin^2 \theta +2\sin\theta \cos \theta +\cos^2 \theta \]

OpenStudy (xapproachesinfinity):

recall that \[\sin^2 \theta +\cos^2 \theta =1\]

OpenStudy (xapproachesinfinity):

\[-2\sin \theta \cos \theta +2\sin \theta \cos \theta =0\]

OpenStudy (xapproachesinfinity):

do you see why it is 2 now?

OpenStudy (xapproachesinfinity):

taking the square is cannot be used here at all you can't just take the square root because you see the square so careful taking the square root is used in equations (ie cancels the square)

OpenStudy (onepieceftw):

Where does the Pythagorean identity some in?

OpenStudy (xapproachesinfinity):

we are using square root in equations because doing something in one side of the equation, we need to do the same thing the other side for the equation to hold otherwise it has no sense maintain the balance of an equation

OpenStudy (xapproachesinfinity):

i'm not sure i understand your question, but i used the Pythagoras identity twice to give us 2 sin^2 theta +cos ^2 theta=1 see my reply above

OpenStudy (onepieceftw):

Still don't get it but thanks anyways. :P

OpenStudy (triciaal):

@onePieceFTW did you understand my response?